# Emergence in graphs with near-extreme constraints

### v2

@@ Abstract.
We consider optimal graphons associated with extreme and near-extreme constraints on the densities of edges and triangles. We prove that the optimizers for near-extreme constraints are unique and multipodal and are perturbations of the previously known unique optimzers for extreme constraints. This proves the existence of infinitely many phases. We determine the podal structures in these phases and prove the existence of phase transitions between them.
1. Results and setting
1.1. New results
We investigate the large scale structures called podes which emerge as node number increases in simple graphs with given edge and triangle densities $(e,t)$. In particular we focus on phase transitions which appear as discontinuities in the pode structure as $(e,t)$ is varied.
The boundary curves in Figure 1 show the extreme accessible values of pairs $(e,t)$ of those densities. Using graphons, the Large Deviation Principle (LDP) of $\mathbb{G}(n,p)$ graphs, and a Boltzmann entropy taken from statistical mechanics [37], one can associate [37] with each extreme-density pair $(e,t)$ a unique optimal graphon, i.e. a graphon with the given densities which optimizes the LDP rate function.
@@ 1.1. New results
In this paper we prove that the unique optimal graphons at the extreme densities extend smoothly into unique optimal graphons in the nearby interior of the triangle, and we fully determine their pode structure. An important technical result of the paper is a new technique for proving the existence of these near-extreme phases, in particular distinct phases emanating from each of the infinitely many scallops at the bottom of Figure 1.
1.2. The background
To prove sharp transitions in large simple graphs we depend heavily on the graph-limit or graphon formalism of Lovász et al [7, 8, 26, 27, 28] (following the work of Aldous [1], Hoover [18] and Kellenberg [20]; see [4, 16] for background). Graphons are an infinite-node idealization of graphs. They do not have edges, but they do have well-defined densities of finite graphs, including edges. Emergence contrasts a large scale with a small scale in finite simple graphs. The smallest scale is that of individual edges, and the large scale is that of dense graphs with many edges on many nodes. Our proofs use graphons in place of large finite graphs, but we justify the distinction. We prove the existence of arbitrarily sharp transitions in asymptotically large finite graphs, based on a dependency between edges produced by triangles. In 2010 Chatterjee and Varadhan proved [13] their LDP for $\mathbb{G}(n,p)$, built on top of the graphon formalism. We introduced in [37] the Boltzmann entropy $\mathbb{B}(e,t)$ as the exponential rate of growth of the number of graphs with densities $(e,t)$. Using the rate function $I(g)$ of the LDP, we proved [37] the representation of $\mathbb{B}(e,t)$ as the maximum of $S(g)=-I(g)$ over all graphons $g$ with edge-triangle constraints $(\varepsilon(g),\tau(g))=(e,t)$. The LDP then implied that if this constrained optimization has a unique graphon solution $g_{e,t}$ (an important proviso!), then all but an exponentially small fraction of large but finite graphs with edge and triangle densities close to $(e,t)$ are well-approximated by $g_{e,t}$ [38]. The latter implies that any true singularites proven at the infinite-node graphon scale represent the approximate phenomenon for exponentially-most large finite constrained graphs.
Completion of Figure 1 in 2012 [35] was a significant achievement by Pikhurko and Razborov. This paper is part of a series based on analyzing phase transitions which emerge at various places in the Razborov triangle.
@@ 1.2. The background
Figure 2. Schematic drawing of a conjecture from 2017 [21], based on computer simulations of optimal graphons associated with the phases of large graphs with edge and triangle constraints.
In the initial papers we developed methods to determine a few phases far in the interior of the Razborov triangle, near the curve $t=e^{3}$ associated with graphs with independent edges. We defined a phase as a connected open set in the $(e,t)$ plane in which the unique optimal graphons, and therefore $\mathbb{B}$, vary smoothly with $(e,t)$, so that phase transitions would be seen as singularities separating phases. This assumed that smoothness prevailed over most of Figure 1 so one could concentrate on the transitions. Although this turned out to be the case, the proof took several years.
Our project thus amounted to showing that, apart from rare (codimension-1) values of $(e,t)$, the $S(g)$-maximizing graphon with densities $(e,t)$ is unique and varies smoothly with $(e,t)$, implying that $\mathbb{B}(e,t)$ is smooth at $(e,t)$. From computer simulations we were led in 2017 to conjecture [21] that this is the case except across the curves shown in Figure 2. Consequently, all unconstrained subgraph densities associated with exponentially-most graphs should be piecewise-smooth functions of $(e,t)$, with singularities at phase boundaries. This conjecture gave the project a more concrete goal.
Progress was slow until now. We first proved that the picture in Figure 2 was correct just above the Erdős-Rényi curve, $t=e^{3}$, (the $F(1,1)$ phase in the figure) [22], then just below the Erdős-Rényi curve when $e>1/2$ (the $B(1,1)$ phase) [32], and then a neighborhood of the $e=1/2$ line for $t<e^{3}$ (part of the $A(2,0)$ phase) [39, 33].
1.3. Connections with other optimization problems
@@
By far the most highly developed mathematical formalism concerned with emergence is equilibrium statistical mechanics. In our project we have sought to create an analogous formalism for simple constrained graphs, motivated by the Razborov triangle, Figure 1. One of our goals is to better understand constrained graphs, to extend the classic study of extremal graphs. Another is to provide an example for other optimization problems facing similar obstacles. For this reason we give a brief sketch of emergence in statistical mechanics [41, 47, 46], emphasizing the significance of convexity.
@@
A common feature of edge-triangle graphs and statistical mechanics is constrained optimization. There are several ways to view equilibrium statistical mechanics as a constrained optimization problem [19, 17, 40, 23], all of which involve the global optimization of any of a range of free energy functionals, or the entropy. The entropy in statistical mechanics is a measure of the number of possible particle configurations with given constraints. It is a fundamental quantity. It is no exageration to view statistical mechanics as built on the convexity of this entropy (see the lectures of Lanford in [23] and the introduction by Wightman in [19]). The convexity of the entropy allows one to analyze the system without loss of information by the use of a variety of free energies [46], such as the Gibbs free energy $G(p,T)$, which is more familiar than the entropy.
@@
Figure 3 shows a primitive thermodynamic phase diagram, illustrating the pattern of solid and fluid phases in a physical bulk material (i.e. the large scale in the emergent picture), as functions of constraint parameters pressure $p$ and temperature $T$ [36]. The experimentally measurable Gibbs free energy $G(p,T)$ [47, 42] is found to vary smoothly within each region, but shows singular behavior when constraints cross some lower dimension curves (see [40] and section VI in [19]), where bulk material properties such as mass density and heat capacity can change abruptly.
@@
In the analogy between constrained graphs and statistical mechanics, discussed for instance in [34], [44] and [12], edges play the role of individual particles and the energy of a graph is the number of copies of some subgraph $H$, such as triangles. A key to understanding phases in both statistical mechanics and constrained graphs is defining an order parameter [3], which is a function that is identically zero in one phase but nonzero in another. We show its use in Section 4.6.
@@
Unfortunately, with graphs there is no analog of equivalent free energies. This is analyzed in Section 6. This was the starting point of our project in 2013 [37]. (The “inequivalence of ensembles” in constrained graphs was rediscovered and analyzed later; see [15].) This inequivalence presents a serious obstacle. In statistical mechanics free energies provide considerable technical advantages. In their absence we had to develop replacement tools. One such tool, which we introduce in this paper, is the “worth” $W(C)$ that we associate to each column $C$ of a graphon. Developing our variational arguments at the level of columns of graphons is a significant advance in our understanding of optimal graphons and therefore phase transitions in constrained graphs.
@@
We note that some important optimization problems suffer from nonconvexity complications. For instance, the form of optimal transport theory developed by Kantorovich and Brenier was based on convex analysis. It required significant developments over several years to allow the original nonconvex Monge problem to access that convex analysis. (For an introduction to optimal transport see the preface in [45] or chapter 1 in [29]).
@@
The emergence of structure in large but finite physical materials is dramatic, particularly the diversity of solid phases, such as the graphite and diamond phases of carbon. All the richness displayed by the phases of all materials, not just the pure elements but also the huge number of compounds like water and alcohol, is created by the electromagnetic interaction within and between molecules. All of inorganic chemistry comes down to the 100 or so different (integer) electric charges of atomic nuclei.
@@
Emergence gives rise to the complicated structure of “water” in Figure 3 and in the edge-triangle model in Figure 2. The study of the emergence of such diversity from the interaction of invisibly small components of a small number of types has led to a great deal of interesting mathematics in the past twenty years. It is this promising history which was the motivation to bring the richness of statistical mechanics, which is built on the convexity of its entropy, into the nonconvex setting of the Boltzmann entropy $\mathbb{B}$ of constrained graphs.
@@ 1.4. Definitions
In order to state our results more precisely, we need a series of definitions. For an in-depth presentation of graphons see [25]. For an introduction to Large Deviation Principles (and also a compact introduction to graphons) see [9].
Here are a few of the basic definitions. A graphon is a measurable function $g:[0,1]^{2}\to[0,1]$, with $g(y,x)=g(x,y)$ for every ordered pair $(x,y)$. Two graphons $g_{1}$ and $g_{2}$ are said to be equivalent if there is a measure-preserving transformation $\sigma:[0,1]\to[0,1]$ such that $g_{2}(x,y)=g_{1}(\sigma(x),\sigma(y))$ for almost every $(x,y)$. The equivalence classes of graphons are called reduced graphons. The Shannon entropy, edge density and triangle density of a graphon are given by the functionals
[…]
@@
is the usual entropy of independent coin flips with probability $u$ of getting heads. Note that if $g_{1}$ and $g_{2}$ are equivalent graphons, then $S(g_{1})=S(g_{2})$, $\varepsilon(g_{1})=\varepsilon(g_{2})$ and $\tau(g_{1})=\tau(g_{2})$. These quantities should really be viewed as functionals on the space of reduced graphons. $S(g)$ is the negative of the LDP rate function $I(g)$.
The following definitions are less standard.
If the graphon $g$ maximizes $S(g)$ subject to constraints on the densities on edges and another graph, we say that $g$ is an optimal graphon at those densities. When we speak of an optimal graphon being unique, we mean that the reduced graphon is unique. In other words, that the optimal graphon is unique up to equivalence.

### v3

@@ Abstract.
We consider entropy-optimal graphons associated with extreme and near-extreme constraints on the densities of edges and triangles. We prove that the optimizers for near-extreme constraints are unique and multipodal and are perturbations of the previously known unique optimzers for extreme constraints. This proves the existence of infinitely many phases. We determine the podal structures in these phases and prove the existence of phase transitions between them.
1. Setting and results
In this paper we determine the structure of typical large graphs with a specified density $e$ of edges and a near-extreme density $t$ of triangles. It is known that the structure of all but an exponentially small fraction of such graphs are described by a graphon that maximizes a certain entropy functional subject to certain integral constraints, assuming that the entropy maximizer is unique up to equivalence. (See Section 1.3 for relevant background.) We show that this graphon is indeed unique and takes a simple multipodal form, with parameters that are piecewise analytic functions of $(e,t)$. We thereby prove the existence of infinitely many phases and infinitely many phase transitions. These results are stated in Theorems 1, 2, 3 and 4.
Our solutions to the entropy maximization problem rely on the theory of Lagrange multipliers and on a new quantity, called “worth”, that applies to columns of graphons. In Section 2 we develop the theory of Lagrange multipliers for graphons, define worth, and prove in Theorem 11 that all of the columns of an entropy-maximizing graphon must maximize worth.
1.1. The background
The boundary curves in Figure 1 show the extreme accessible values of pairs $(\varepsilon,\tau)$ of densities of edge and triangle subgraphs in asymptotically large simple graphs [45]. Values throughout the interior are also accessible, and can be studied, using the formalism of graph limits or graphons developed in 2006-2011 by Lovász et al [7, 8, 30, 31, 32] (see [4, 17] and [29] for background), and the Large Deviation Principle (LDP) of $\mathbb{G}(n,p)$ graphs [14].
@@ 1.1. The background
The structure of the graphs on the boundary of Figure 1 appeared in 2012 in a preprint of [40]. Following as it did soon after the development of graphons and the LDP for $\mathbb{G}(n,p)$, this led to a series of papers aimed at analyzing emergent (large scale) structures when the edge/triangle constraints $(\varepsilon,\tau)$ were near the “Erdős-Rényi” curve $\tau=\varepsilon^{3}$, associated with graphs in $\mathbb{G}(n,p)$. These papers used the LDP to introduce a Boltzmann entropy, $\mathbb{B}(e,t)$, which measures the exponential rate of growth of the number of large graphs with edge and triangle densities $(\varepsilon,\tau)\approx(e,t)$, a notion copied from statistical mechanics. See [43] for a precise statement and proof of this theorem, and the important fact that $\mathbb{B}(e,t)$ is not convex in $(e,t)$.
Computer simulations led to the conjecture [26] of many distinct emergent “phases” throughout the interior of Figure 1, as seen in Figure 2. Eventually emergent phases were determined for parts of the three regions $F(1,1)$ [25], $B(1,1)$ [36] and $A(2,0)$ [37] in Figure 2 near the Erdős-Rényi curve.
@@
Figure 2. Schematic drawing of a conjecture from 2017 [26], based on computer simulations of entropy-optimal graphons associated with the phases of large graphs with edge and triangle constraints.
1.2. Connections with other research
There is widespread interest in analyzing network and graph data [38], and statistical methods have developed to support this, including a family of Exponential Random Graph Models (ERGMs). In [13] a number of practical problems with ERGMs were isolated and treated using the large deviations principal of Erdős-Rényi graphs. In [42] a Boltzmann entropy $\mathbb{B}(\bar{\tau})$ was introduced which, together with the LDP, allowed the analysis of ‘exponentially most’ large finite graphs with constraints on the densities $\bar{\tau}$ of some subgraphs; see [43] for a review of these results in a general setting. Using $\mathbb{B}(\bar{\tau})$ some of the weaknesses of ERGMs discussed in [13] can be quantified. This is discussed in detail in Section 6.
@@ 1.2. Connections with other research
Our paper is concerned with the emergence [28] of large scale structure in graphs, for instance of podes. By far the most highly developed mathematical formalism concerned with emergence is equilibrium statistical mechanics. In our work we have sought to create an analogous formalism for simple constrained graphs, motivated by the Razborov triangle, Figure 1.
One of our goals is to better understand constrained graphs, to extend the classic study of extremal graphs, of which [40] is a distinguished result [2, 6]. Another is to provide an example for other optimization problems facing similar obstacles. For this reason we give here a brief sketch of emergence in statistical mechanics [47, 53, 52], emphasizing the significance of entropy, free energies and especially convexity.
@@
A common structure underlying the mathematics of edge-triangle graphs and the mathematics of statistical mechanics is constrained optimization. There are several ways to view equilibrium statistical mechanics as constrained optimization on a space of many-particle configurations [23, 19, 46, 27], all of which involve the global optimization of any of a range of free energy functionals, or the entropy. The entropy in statistical mechanics is a measure of the number of possible particle configurations with given constraints. It is a fundamental quantity. It is no exageration to view statistical mechanics as built on the convexity of this entropy (see the lectures of Lanford in [27] and the introduction by Wightman in [23]).
The convexity of the entropy allows one to analyze the system without loss of information by the use of a variety of free energies [52], such as the Gibbs free energy $G(p,T)$, which is more familiar than the entropy.
@@
Figure 3 shows a primitive thermodynamic phase diagram, illustrating the pattern of solid and fluid phases in a physical bulk material (the phases represent the large particle-number scale in the emergent picture), as functions of constraint parameters pressure $p$ and temperature $T$ [41]. The experimentally measurable Gibbs free energy $G(p,T)$ [53, 48] is found to vary smoothly within each region, but shows singular behavior when constraints cross some lower dimension curves (see [46] and section VI in [23]), where bulk material properties such as mass density and heat capacity can change abruptly.
@@
In the analogy between constrained graphs and statistical mechanics, discussed for instance in [39], [50] and [13], a large simple graph $G$ contains many edges, which play the role of individual particles, and the number of copies of some subgraph $H$ in $G$, such as triangles, plays the role of the total (potential) energy of $G$. A key to understanding emergent phases in both statistical mechanics and constrained graphs is defining an order parameter [3], which is a function that is identically zero in one phase but nonzero in another. We show its use in Section 4.6.
@@
Unfortunately, with graphs there is no analog of equivalent free energies. (This is analyzed in Section 6.) This ‘inequivalence of ensembles’ presents a serious obstacle. In statistical mechanics free energies provide considerable technical advantages. In their absence we had to develop replacement tools. One such tool, which we introduce in this paper, is the “worth” $W(C)$ that we associate to each column $C$ of a graphon.
@@
We note that some important optimization problems also suffer from nonconvexity complications. For instance, the form of optimal transport theory developed by Kantorovich and Brenier was based on convex analysis. It required significant developments over several years to allow the original nonconvex Monge problem to access that convex analysis. (For an introduction to optimal transport see the preface in [51] or chapter 1 in [33]).
@@
The emergence of structure in large but finite physical materials is dramatic, particularly the diversity of solid phases, such as the graphite and diamond phases of carbon. (There are more than 20 distinct solid phases of water.) All the richness displayed by the phases of all materials, not just the pure elements but also the huge number of compounds like water and alcohol, is created by the electromagnetic interaction within and between molecules. All of inorganic chemistry comes down to the 100 or so different (integer) electric charges of atomic nuclei.
@@
Emergence gives rise to the complicated structure of “water” in Figure 3 and in the edge-triangle model in Figure 2. The study of the emergence of such diversity from the interaction of invisibly small components of a small number of types has led to a great deal of interesting mathematics in the past twenty years. It is this promising history which was the motivation to bring the richness of emergent structure in statistical mechanics, which is built on the convexity of its entropy, into the nonconvex setting of the Boltzmann entropy $\mathbb{B}$ of constrained graphs.
@@ 1.3. Definitions
We include here the definitions and concepts needed to state our results precisely and to prepare the reader to follow the proofs.
1.3.1. Graphons and the cut topology
asdf
Intuitively, graphons are limits, as node number $n\to\infty$, of the adjacency matrices of simple graphs, with the nodes mapped to the interval $[0,1]$. For background up to its publication in 2011 and an encylopedic treatment of various aspects of graphons we recommend the book [29] by Lovász. For a more recent treatment, see [10].
Definition of graphons. The space $\mathcal{W}$ of graphons is the quotient, of the space of Borel measurable functions $g:[0,1]^{2}\to[0,1]$ satisfying $g(x,y)=g(y,x)$, by the identification of functions that differ only on a set of Lebesgue measure zero.
The cut topology on $\mathcal{W}$ is metrizable with the following metric.
Definition of the cut metric.
(1) $d_{cut}(g_{1},g_{2})\equiv\sup_{S,T}\Big{|}\int_{S\times T}\big{[}g_{1}(x,y)-g_{2}(x,y)\big{]}\,dx\,dy\Big{|},$
where $S,T$ are measurable subsets of $[0,1]$.
Note that the absolute value goes outside the integral!
1.3.2. Subgraph densities and weak equivalence
asdf
For a simple graph $F=(V,E)$ on $k$ nodes, we define the (homomorphism) density of $F$ in the graphon $g$ as
(2) $\tau_{F}(g)=\int_{[0,1]^{k}}\prod_{{i,j}\in E}g(x_{i},x_{j})dx_{1}\,dx_{2}\,\ldots,dx_{k}.$
It can be proven [10] that $\tau(F)$ is continuous on $\mathcal{W}$ in its cut topology. We will concentrate on the densities
(3) $\varepsilon(g)=\iint g(x,y)\,dx\,dy$
and
(4) $\tau(g)=\iiint g(x,y)g(y,z)g(z,x)\,dx\,dy\,dz$
of edges and triangles, respectively.
The space $\mathcal{W}$ of ordinary graphons is not compact in the cut topology. To obtain a compact space $\widetilde{\mathcal{W}}$ we take the quotient of $\mathcal{W}$ by a weak equivalence relation associated with subgraph densities.
Definitions of weak equivalence and reduced graphons. Graphons $g_{1}$ and $g_{2}$ are weakly equivalent if $\tau_{F}(g_{1})=\tau_{F}(g_{2})$ for every simple subgraph $F$. Elements of the quotient space $\widetilde{\mathcal{W}}$ are called reduced graphons.
For our purposes, it is useful to use a different description of weak equivalence. The group of measure-preserving transformations of $[0,1]$ acts naturally on $\mathcal{W}$. If $g$ is a graphon and $\sigma$ is such a transformation, we define, in terms of any representative modulo measure zero:
(5) $g^{\sigma}(x,y)=g(\sigma(x),\sigma(y)).$
We say that $g$ and $g^{\sigma}$ are group equivalent. Note that $\tau_{F}(g^{\sigma})=\tau_{F}(g)$, thanks to a simple change-of-variables in the integral (2), so group equivalence implies weak equivalence.
We next define a pseudometric $\delta_{cut}$ on $\mathcal{W}$ that measures how far two graphons are from being group equivalent.
Definition of $\delta_{cut}$.
(6) $\delta_{cut}(g_{1},g_{2})=\inf_{\sigma_{1},\sigma_{2}}d_{cut}(g_{1}^{\sigma_{1}},g_{2}^{\sigma_{2}}).$
It can be proven [10] that $\delta_{cut}(g_{1},g_{2})=0$ if and only if $g_{1}$ and $g_{2}$ are weakly equivalent. The pseudometric $\delta_{cut}$ on $\mathcal{W}$ then descends to a metric $\delta_{cut}$ on $\widetilde{\mathcal{W}}$: $\delta_{cut}([g_{1}],[g_{2}])=\delta_{cut}(g_{1},g_{2})$, where $[g]$ denotes the weak equivalence class of $g\in\mathcal{W}$.
1.3.3. Compactness, the LDP of $\mathbb{G}(n,p)$ graphs and the Shannon entropy
asdf
A major result in the graphon formalism is the
Compactness theorem. $\widetilde{\mathcal{W}}$ is compact in the topology defined by the metric $\delta_{cut}$.
Our results rely heavily on the LDP of $\mathbb{G}(n,p)$ graphs, which is expressed in terms of graphons. For an elegant and concise reference for both the graphon formalism and the LDP we recommend the book [10] of Chatterjee. [^1] Note, however, that Chatterjee works with a slightly different metric on $\mathcal{W}$, defining $d^{\prime}_{cut}(g_{1},g_{2})=\sup_{a,b}\Big{|}\int_{[0,1]^{2}}a(x)b(y)\big{[}g_{1}(x,y)-g_{2}(x,y)\big{]}\,dx\,dy\Big{|},$ where $a$ and $b$ are Borel measurable maps from $[0,1]$ to [1,1][-1,1]. $d^{\prime}_{cut}(g_{1},g_{2})$ is bounded above and below by multiples of $d_{cut}(g_{1},g_{2})$, so Chatterjee’s topological results based on $d^{\prime}_{cut}$ and the corresponding $\delta^{\prime}_{cut}$ apply equally well to $d_{cut}$ and $\delta_{cut}$.
We next introduce some terms associated with the LDP.
The Shannon entropy of a graphon $g$ is
(7) $S(g)=\iint H\big{(}g(x,y)\big{)}\,dx\,dy,$
@@ 1.3.3. Compactness, the LDP of $\mathbb{G}(n,p)$ graphs and the Shannon entropy
is the usual entropy of independent coin flips with probability $u$ of getting heads. Note that $H(u)$ is concave down and that $H^{\prime}(u)=\ln\left(\frac{1-u}{u}\right)$ diverges as $u$ approaches 0 or 1. $S(g)$ is invariant under measure-preserving transformations of $[0,1]$ and defines a function (also denoted $S$) on $\widetilde{\mathcal{W}}$. $S$ is also minus the rate function of the LDP of $\mathbb{G}(n,p)$ graphs. The LDP relates the number of large graphs associated with an (open or closed) subset of $\widetilde{\mathcal{W}}$ with the supremum of $S$ on that subset; see [10] for more details on the LDP.
Let $\mathcal{W}_{e,t}$ be the set of graphons with $\varepsilon(g)=e$ and $\tau(g)=t$ and let $\widetilde{\mathcal{W}}_{e,t}$ be the corresponding set of reduced graphons.
The Boltzmann entropy function $\mathbb{B}(e,t)$ can be understood in two ways. One is as the exponential rate of growth, as the number of nodes $n$ diverges, of the number of graphs on $n$ nodes with edge/triangle densities $(e,t)$. For this paper, it is useful to use the fact, proven in [42, 44], that $\mathbb{B}(e,t)$ equals the maximum of $S([g])$ on $\widetilde{\mathcal{W}}_{e,t}$, which is the same as the maximum of $S(g)$ on $\mathcal{W}_{e,t}$.
Most of this paper is devoted to maximizing $S$ on $\mathcal{W}_{e,t}$ or $\widetilde{\mathcal{W}}_{e,t}$. That is, we focus on the constrained optimization of $S$. The optimizing graphons are called entropy-optimal graphons (or “optimal graphons”, for short) for the given constraints $(e,t)$. When we refer to an optimal graphon being unique we always mean that the optimal reduced graphon is unique.
The conjecture [44, 26] that constrained optimization of $S$ would give rise to the rich phenomena of Figure 2, much of which is finally proven in this paper, was by analogy of $\mathbb{B}(e,t)$ with the Boltzmann entropy in statistical mechanics, as discussed in subsection 1.2.
1.3.4. Multipodality, phases and phase transitions
asdf